# SSL-OTA: Unveiling Backdoor Threats in Self-Supervised Learning for Object Detection

### v1

@@ Abstract
The extensive adoption of Self-supervised learning (SSL) has led to an increased security threat from backdoor attacks. While existing research has mainly focused on backdoor attacks in image classification, there has been limited exploration into their implications for object detection. In this work, we propose the first backdoor attack designed for object detection tasks in SSL scenarios, termed Object Transform Attack (SSL-OTA). SSL-OTA employs a trigger capable of altering predictions of the target object to the desired category, encompassing two attacks: Data Poisoning Attack (NA) and Dual-Source Blending Attack (DSBA). NA conducts data poisoning during downstream fine-tuning of the object detector, while DSBA additionally injects backdoors into the pre-trained encoder. We establish appropriate metrics and conduct extensive experiments on benchmark datasets, demonstrating the effectiveness and utility of our proposed attack. Notably, both NA and DSBA achieve high attack success rates (ASR) at extremely low poisoning rates (0.5%). The results underscore the importance of considering backdoor threats in SSL-based object detection and contribute a novel perspective to the field.
Index Terms—  Backdoor Attack, Self-Supervised Learning, Object Detection
1 Introduction
@@
In the SSL domain, there has been research on backdoor attacks in image classification [7, 8, 9]. However, such attacks on object detection remain unexplored, with limited work in supervised learning  [10, 11, 12, 13]. Wu et al. [13] attempted to use object rotation as a backdoor trigger, but this was circumvented by data augmentation in SSL, suggesting SSL’s potential in enhancing adversarial robustness in object detection models. Compared to image classification, object detection is widely applied in critical tasks such as pedestrian detection [14] and autonomous driving [15]. Backdoor attacks on object detection models may pose a more serious threat to human life and property. For instance, a potential backdoor trigger that prevents the model from recognizing a person could lead to serious traffic accidents. Attacking object detectors is more challenging than classifiers because it involves classifying and locating multiple targets within an image, and object detection models, like Fast R-CNN [16], are more complex than image classification models. Therefore, in this study, we investigate how to design effective backdoor attacks specifically tailored for object detection tasks.
@@
In this paper, we propose the first backdoor attacks on object detection in the context of SSL, termed Object Transform Attack (SSL-OTA). SSL-OTA employs a trigger to alter the category of the targeted object to the desired class. Subsequently, we delineate two attacks for OTA, namely NA and DSBA. NA involves poisoning the training set with trigger-containing toxic samples during the downstream fine-tuning phase and alters the ground truth labels (object categories) of the poisoned images. In this scenario, the Attackers have neither information nor the ability to control other training components. However, the attack capability is limited. To enhance the attack effectiveness of the attack, we further introduce DSBA, a hybrid attack leveraging dual data sources from both the encoder and the downstream detector. This involves backdoor injection into the pre-trained encoder using an additional shadow dataset, while maintaining the same settings as NA in the downstream phase.
@@
To evaluate the effectiveness of our attacks, we design appropriate evaluation metrics, including mAP and AP computed on the backdoored test dataset and benign test dataset. We conduct extensive experiments on two benchmark datasets and two widely used model architectures. Empirical results indicate that both attacks achieve high ASR while maintaining model utility. For instance, when targeting the downstream object detector constructed on PASCAL VOC2007 [17], NA demonstrates an ASR of 72.56%, while DSBA achieves an ASR of 86.55%. Furthermore, Furthermore, in most cases, the utility loss inflicted by both attacks is within 1%. We also conduct comprehensive ablation studies to analyze factors influencing attack performance, demonstrating the robustness of our backdoor attacks across different settings. Our key contributions are summarized as follows:
@@
  • Unveiling backdoor threats in the SSL scenario targeted at object detection. To our knowledge, this marks the first instance of a backdoor attack on this critical task.
@@
  • Based on the characteristics of target detection, we designed two simple yet effective stealthy backdoor attacks, NA and DSBA.
@@
  • We validated the effectiveness and utility of our attacks through extensive experiments on benchmark datasets.
2 Threat Model
@@
Attacker’s Capacities. In the attack scenarios delineated, we assume for both attacks that the adversary can inject a minimal amount of data samples into the training dataset. Additionally, the DSBA considers two potential attackers: 1) untrusted service providers who inject backdoors into their pre-trained image encoders and offer them to users, and 2) malicious third parties who insert backdoors into service providers’ pre-trained image encoders and release them online for user download. DSBA also assumes that attackers possess a shadow dataset with a distribution similar to the downstream dataset. Notably, for both attacks, the attacker cannot access the downstream dataset or manipulate the training process of the downstream classifier.
Attacker’s Goals. The attacker has two main targets, including 1) Effectiveness, ensuring that when the specified trigger pattern by the attacker is present on an object of a designated category, the compromised model will misclassify this object as the target category. 2) Utility, requiring the compromised model’s ability to detect benign objects with performance comparable to a model trained on a benign training dataset.
3 Methodology
@@
Fig. 1: The main pipeline of NA and DSBA against object detection. Our method involves two attacks of SSL-OTA: NA and DSBA. NA poisons only a small number of training samples during the downstream fine-tuning phase(right side). In contrast, DSBA conducts a hybrid attack using dual data sources from the encoder and downstream detector(left side). Building upon NA, it introduces additional shadow dataset for backdoor injection into the shadow object detector, comprising three steps: 1) Shadow object detector training, 2) Backdoor training and extraction, and 3) Poisoned fine-tuning. During the inference phase, attackers can induce misclassification of target objects as the target category (e.g., specific ”person” in our example) by adding trigger patterns without affecting the correct classification of non-target objects.
3.1 Object Transform Attack
Definition. The function $G(\cdot)$ represents the trigger generator, and $M_{infected}$ is the infected model. For any given image $\boldsymbol{x}$, where $\boldsymbol{x}$ contains multiple detection objects $o_{i}$(where i=1,2,…,n), each object belonging to category $c_{i}$, SSL-OTA satisfies the following conditions: $M_{\text{infected }}\left(x\left(G\left(o_{i}\right),o_{j}\right)\right)=\left(c_{t},c_{j}\right)$, where $c_{t}$ represents the target class.
3.2 Naive Attack
@@
$\mathcal{D}_{u}=\left\{\left(G_{i}(\boldsymbol{x}),G_{a}(\boldsymbol{a})\right)\mid(\boldsymbol{x},\boldsymbol{a})\in\mathcal{D}_{l}\right\}$ (1)
@@
where $G_{i}$ and $G_{a}$ represent the poisoned image generator and the poisoned annotation generator, respectively. The trigger injection formula for $G_{i}$ is:
$G_{i}(\boldsymbol{x})=(\boldsymbol{1}-\boldsymbol{\alpha})\otimes\boldsymbol{x}+\boldsymbol{\alpha}\otimes\boldsymbol{t}$ (3)
where $t$ represents the trigger pattern specified by the adversary, and $\boldsymbol{\alpha}\in[0,1]^{C\times W\times H}$ denotes the trigger’s transparency. The poisoned annotation generator $G_{a}$ alters the true label of the target objects to the target label $c_{t}$.
@@ 3.3 DSBA
DSBA is an enhanced attack built upon NA, considering the perspective of a SSL pre-trained encoder, employing a blended attack strategy that utilizes both the encoder and downstream detector as dual data sources. The attack process is as follows:
Shadow Object Detector Training. We observed that directly loading the pre-trained encoder for backdoor training would compromise the independence of the encoder’s backdoor, especially when using other detection structures without normal performance. Therefore, we initially employ a SSL pre-trained encoder to construct a clean object detector downstream, serving as the shadow object detector. Let $\mathcal{D}=\left\{\left(\boldsymbol{x}_{i},\boldsymbol{a}_{i}\right)\right\}_{i=1}^{N}$ denote the benign dataset,where $x_{i}\in X$ represents the image of an object,, and $a_{i}\in A$ is the ground-truth annotation of the object $x_{i}$. For each annotation $a_{i}$, we have $a_{i}=[x_{i},y_{i},w_{i},h_{i},c_{i}$, where $(x_{i},y_{i})$ is the center coordinates of the object, $w_{i}$ is the width of the bounding box(bbox), $h_{i}$ is the height of the bbox, and $c_{i}$ is the class of object $x_{i}$. Given the dataset $D_{clean}$, the user trains the object detector through the standard training process.
@@
Backdoor Training and Extraction. We performed backdoor injection into the encoder component of the shadow object detector, using an additional shadow dataset distinct from the downstream dataset, while keeping the remaining parts frozen. The method for generating the backdoor data is consistent with the previous description in NA. Specifically, we added a trigger pattern to all target bboxes of the attacked class in the images of the benign dataset $D_{clean}$. Subsequently, we used the poisoned annotation generator $G_{a}$ to modify the true labels in their corresponding annotations to the target label $c_{t}$, resulting in the shadow dataset $D_{shadow}$:
$\mathcal{D}_{shadow}=\left\{\left(G_{i}(\boldsymbol{x}),G_{a}(\boldsymbol{a})\right)\mid(\boldsymbol{x},\boldsymbol{a})\in\mathcal{D}_{shadow}\right\}$ (4)
Then we conducted backdoor training using the shadow dataset $D_{shadow}$, i.e.:
$\min\frac{1}{N}\sum_{(\boldsymbol{x},a)\in\mathcal{D}_{shadow}}\mathcal{L}\left(f(\boldsymbol{x}),a\right)$ (5)
where $\mathcal{L}$ is the loss function.
The attacker then extracts the backdoor encoder from the shadow object detector obtained through backdoor training, and releases it.
Poisoning Fine-tuning. The attacker generates a poisoned dataset and releases it. Notably, the poisoning ratio in the poisoned dataset is extremely low, such as 0.5%, making the attack more closely resemble real-world scenarios. Upon receiving the backdoor encoder and the poisoned dataset, the victim conducts fine-tuning for their downstream task. The fine-tuning process aligns with the procedure described in NA.
4 Evaluation
4.1 Experiment settings
Datasets. In this work, we utilized the PASCAL VOC2007 [17], PASCAL VOC07+12 [18] and MSCOCO [19] datasets for a comprehensive evaluation. Each image is annotated with bboxes coordinates and corresponding classes. The VOC2007 dataset comprises 20 categories, divided into training/validation data (trainval) and testing data (test). The MSCOCO trainval dataset encompasses 80 categories, with 118k training images utilized for training (58k as the fine-tuning dataset and 60k as the shadow dataset), while the validation set serves as the test data. Throughout all experiments, we selected the ”person” category as the attacked class and the ”dog” category as the target class.
@@ 4.1 Experiment settings
Table 1: Attack performance of the two attacks on object detection. Note that $\uparrow$”/$\downarrow$”/“-” indicate the metric should be high/low/similar to the same metric of $M_{benign}$ to show the success of the attack.
Downstream Dataset Model Method $mAP_{benign}$ $AP_{benign}$ $mAP_{attack}$uparrow $AP_{attack}$uparrow $mAP_{attack+benign}$ $AP_{attack+benign}$downarrow $-mAP_{part\_attack+benign}\text{-}$ ASR $\uparrow$
VOC2007 Fast R-CNN NAIVE ATTACK 68.69 70.26 62.77 80.01 60.47 1.46 58.98 72.56
DSBA 70.75 68.19 66.25 89.48 62.84 1.54 60.54 86.55
MSCOCO Fast R-CNN NAIVE ATTACK 51.76 50.54 48.64 45.54 47.19 0.87 46.12 69.95
DSBA 53.03 52.58 50.27 55.98 48.72 0.0 47.54 70.33
Mask R-CNN NAIVE ATTACK 53.45 60.45 50.26 49.52 48.79 0.0 44.77 70.51
DSBA 53.97 55.57 51.04 56.44 49.47 0.0 48.42 71.63

### v2

@@ Abstract
The extensive adoption of Self-supervised learning (SSL) has led to an increased security threat from backdoor attacks. While existing research has mainly focused on backdoor attacks in image classification, there has been limited exploration of their implications for object detection. Object detection plays a critical role in security-sensitive applications, such as autonomous driving, where backdoor attacks seriously threaten human life and property. In this work, we propose the first backdoor attack designed for object detection tasks in SSL scenarios, called Object Transform Attack (SSL-OTA). SSL-OTA employs a trigger capable of altering predictions of the target object to the desired category, encompassing two attacks: Naive Attack (NA) and Dual-Source Blending Attack (DSBA). NA conducts data poisoning during downstream fine-tuning of the object detector, while DSBA additionally injects backdoors into the pre-trained encoder. We establish appropriate metrics and conduct extensive experiments on benchmark datasets, demonstrating the effectiveness of our proposed attack and its resistance to potential defenses. Notably, both NA and DSBA achieve high attack success rates (ASR) at extremely low poisoning rates (0.5%). The results underscore the importance of considering backdoor threats in SSL-based object detection and contribute a novel perspective to the field.
Index Terms:
Backdoor Attack, Self-Supervised Learning, Object Detection
I Introduction
@@
Figure 1: Illustration of the proposed OTA on object detection. This involves each trigger causing the model to misclassify an object of the attacked class (in this case, “person”) as the target class “dog”. We show the predicted bounding boxes with a confidence score $>$ 0.5.
@@
In the SSL domain, there has been research on backdoor attacks in image classification [7, 8, 9, 10]. However, such attacks on object detection remain unexplored, with limited work in supervised learning  [11, 12, 13, 14]. Wu et al. [14] attempted to use object rotation as a backdoor trigger, but this was circumvented by data augmentation in SSL, suggesting SSL’s potential to enhance adversarial robustness in object detection models. Compared to image classification, object detection is widely applied in critical tasks such as pedestrian detection [15] and autonomous driving [16]. Backdoor attacks on object detection models may pose a more serious threat to human life and property. For instance, a potential backdoor trigger that prevents the model from recognizing a person could lead to serious traffic accidents. Attacking object detectors is more challenging than classifiers because it involves classifying and locating multiple targets within an image, and object detection models, like Fast R-CNN [17], are more complex than image classification models. Additionally, backdoor attacks in image classification typically aim to misclassify images into a predetermined target category. However, this approach is unsuitable for object detection tasks. In object detection, a single image contains multiple objects, each with distinct categories and locations. Therefore, in this study, we investigate how to design effective backdoor attacks specifically tailored for object detection tasks.
In this paper, we propose the first backdoor attack on object detection in the context of SSL, termed Object Transform Attack (SSL-OTA). SSL-OTA employs a trigger to alter the category of the targeted object to the desired class. Figure 1 provides an example of the setup. Multiple triggers are inserted into the image, and the compromised model is expected to detect and misclassify the attacked object (“person”) in the poisoned image as the target category (“dog”). In real-world scenarios, such as smart access control systems, misclassifying an authorized person as a dog could lead to unauthorized access or incorrect recording of personal information, thereby posing security risks.
@@
Subsequently, we delineate two types of SSL-OTA, namely Naive Attack (NA) and Dual-Source Blending Attack (DSBA). NA involves poisoning the training set with trigger-containing toxic samples during the downstream fine-tuning phase and alters the ground truth labels (object categories) of the poisoned images. In this scenario, the Attackers have neither information nor the ability to control other training components. However, the attack capability is limited. To enhance the attack effectiveness of the attack, we further introduce DSBA, a hybrid attack leveraging dual data sources from both the encoder and the downstream detector. This involves backdoor injection into the pre-trained encoder using an additional shadow dataset, while maintaining the same settings as NA in the downstream phase.
@@
To evaluate the effectiveness of our attacks, we design appropriate evaluation metrics, including mAP and AP computed on the backdoored test dataset and benign test dataset. We conduct extensive experiments on two benchmark datasets and two widely used model architectures. Empirical results indicate that both attacks achieve high ASR while maintaining model utility. For instance, when targeting the downstream object detector constructed on PASCAL VOC2007 [18], NA demonstrates an ASR of 72.56%, while DSBA achieves an ASR of 86.55%. Furthermore, in most cases, the utility loss inflicted by both attacks is within 1%. We also conduct comprehensive ablation studies to analyze factors influencing attack performance, demonstrating the robustness of our backdoor attacks across different settings. Our key contributions are summarized as follows:
@@
  • Unveiling backdoor threats in the SSL scenario targeted at object detection. To our knowledge, this marks the first attempt of backdoor attacks on this critical task.
@@
  • Based on the characteristics of object detection, we design two simple yet effective stealthy backdoor attacks, NA and DSBA.
@@
  • We evaluate the effectiveness and utility of our attacks through extensive experiments on benchmark datasets.
II Background
II-A Self-supervised learning
SSL has made significant advancements in the field of computer vision in recent years. The core idea is to use a large amount of unlabeled data for pre-training, thereby reducing reliance on labeled data and significantly enhancing the model’s generalization capability. The SSL pipeline consists of two key components: pre-training an image encoder and constructing an object detector (depending on the downstream task). Consequently, backdoor attacks on SSL aim to compromise any part of this pipeline. Next, we discuss these two components.
Pre-training an Image Encoder. The first component aims to pre-train the encoder using a large amount of unlabeled data. Among the many methods [2, 1, 19] for pre-training encoders with unlabeled images, contrastive learning is an effective and commonly used approach. The basic principle of contrastive learning is to maximize the similarity between positive pairs (different augmented versions of the same input image) and minimize the similarity between negative pairs (different input images). This approach learns a representation space where similar images generate similar feature vectors, while different images generate distinct feature vectors.
Constructing a Downstream Object Detector. The second component employs the pre-trained encoder as a feature extractor to develop object detectors for downstream tasks, which require minimal or no labeled training data. Specifically, we have multiple labeled training examples for these downstream tasks, referred to as the downstream dataset. The pre-trained image encoder is utilized to produce a feature vector for each image in this dataset, followed by training an object detector using standard supervised learning techniques. For a given test input, the pre-trained image encoder first generates its feature vector, after which the trained object detector identifies the detection box and predicts the corresponding label.
II-B Threat Model
@@
Attacker’s Capacities. In the attack scenarios delineated, we assume for both attacks that the adversary can inject a minimal amount of data samples into the training dataset like prior work [11]. Additionally, the DSBA considers two potential attackers: 1) untrusted service providers who inject backdoors into their pre-trained image encoders and offer them to users, and 2) malicious third parties who insert backdoors into service providers’ pre-trained image encoders and release them online for user download. DSBA also assumes that attackers possess a shadow dataset with a distribution similar to the downstream dataset. Notably, for both attacks, the attacker cannot access the downstream dataset or manipulate the training process of the downstream classifier.
Attacker’s Goals. The attacker has two main targets, including 1) Effectiveness, ensuring that when the specified trigger pattern by the attacker is present on an object of a designated category, the compromised model will misclassify this object as the target category. 2) Utility, this requires the compromised model’s ability to detect benign objects with performance comparable to a model trained on a benign training dataset.
III Methodology
@@
Figure 2: The main pipeline of NA and DSBA against object detection. Our method involves two attacks of SSL-OTA: NA and DSBA. NA poisons only a small number of training samples during the downstream fine-tuning phase (right side). In contrast, DSBA conducts a hybrid attack using dual data sources from the encoder and downstream detector (left side). Building upon NA, it introduces an additional shadow dataset for backdoor injection into the shadow object detector, comprising three steps: 1) Shadow object detector training, 2) Backdoor training and extraction, and 3) Poisoned fine-tuning. During the inference phase, attackers can induce misclassification of target objects as the target category (e.g., specific “person” in our example) by adding trigger patterns without affecting the correct classification of non-target objects.
III-A Object Transform Attack
In object detection, the goal is to classify and locate objects within an image. Each candidate object outputs a rectangular bounding box (referred to as “bbox”) and a confidence score (ranging from 0 to 1, with higher scores indicating higher confidence). For any given image $\boldsymbol{x}$, which contains multiple detected objects $\boldsymbol{o}_{i}$ (where $i=1,2,...,n$), $\boldsymbol{y}=[\boldsymbol{o}_{1},\boldsymbol{o}_{2},...,\boldsymbol{o}_{n}]$ represents the ground truth labels of $\boldsymbol{x}$. For the bbox $\boldsymbol{o}_{i}$ of a selected category, $\boldsymbol{o}_{i}=[a_{i},b_{i},w_{i},h_{i},c_{i}]$, where $(a_{i},b_{i})$ are the center coordinates of $\boldsymbol{o}_{i}$, $w_{i}$ is the width of the bbox, $h_{i}$ is the height of the bbox, and $c_{i}$ is the category label of $\boldsymbol{o}_{i}$.
Definition. In OTA, $M_{infected}$ is the infected model. For any image $\boldsymbol{x}$ containing multiple detection objects $\boldsymbol{o}_{i}$, we insert a trigger $\boldsymbol{x}_{trigger}$ into the center $(a_{i},b_{i})$ of the bbox $\boldsymbol{o}_{i}$. In this way, we insert $m$ triggers into the image (where $m$ is the number of objects in the selected category), resulting in $\boldsymbol{x}_{poisoned}$. $M_{infected}$ should detect and classify the selected category objects in the $\boldsymbol{x}_{poisoned}$ image as the target category $c_{t}$. The model’s prediction should be $\boldsymbol{y}_{target}=[\boldsymbol{o}_{1},...,\boldsymbol{o}_{n-m},\boldsymbol{o}_{target_{1}},...,\boldsymbol{o}_{target_{m}}]$, where $\boldsymbol{o}_{target_{i}}=[a_{i},b_{i},w_{i},h_{i},c_{t}]$, for $1\leq i\leq m$. Therefore, $M_{infected}(\boldsymbol{x}_{poisoned})=\boldsymbol{y}_{target}$.
III-B Naive Attack
@@
$\mathcal{D}_{u}=\left\{\left(\boldsymbol{x}_{poisoned},\boldsymbol{y}_{target}\right)\mid(\boldsymbol{x},\boldsymbol{y})\in\mathcal{D}_{l}\right\}$ (1)
@@
where $\boldsymbol{x}_{poisoned}$ and $\boldsymbol{y}_{target}$ represent the poisoned image and the poisoned annotation, respectively. The injection formula of the trigger is:
$\boldsymbol{x}_{poisoned}=(\boldsymbol{1}-\boldsymbol{\alpha})\otimes\boldsymbol{x}+\boldsymbol{\alpha}\otimes\boldsymbol{t}$ (3)
where $\boldsymbol{t}$ represents the trigger pattern specified by the adversary, and $\boldsymbol{\alpha}\in[0,1]^{C\times W\times H}$ denotes the trigger’s transparency. In addition, we change the real label of the target object in the poisoning annotation $\boldsymbol{y}_{target}$ to the target label $c_{t}$.
@@ III-C Dual-Source Blending Attack
DSBA is an enhanced attack built upon NA, considering the perspective of an SSL pre-trained encoder, employing a blended attack strategy that utilizes both the encoder and downstream detector as dual data sources. The attack process is as follows:
Shadow Object Detector Training. We found that directly loading a pre-trained encoder for backdoor training, without a properly functioning detection structure, affects the independence of the encoder’s backdoors. Therefore, we first use an SSL pre-trained encoder to construct a clean object detector downstream, ensuring normal performance, which serves as a shadow object detector.
Given a benign dataset $D_{clean}$, let $\mathcal{D}_{clean}=\left\{\left(\boldsymbol{x}_{i},\boldsymbol{y}_{i}\right)\right\}_{i=1}^{N}$, where $\boldsymbol{o}_{i}\in\boldsymbol{x}$ is the target object in the image, and $\boldsymbol{y}=[\boldsymbol{o}_{1},\boldsymbol{o}_{2},\ldots,\boldsymbol{o}_{n}]$ is the ground truth label of $\boldsymbol{x}$. We do not modify the dataset. Instead, we directly use $D_{clean}$ for iterative training, employing a standard training process. Ultimately, we obtain a well-trained shadow object detector.
@@
Backdoor Training and Extraction. We perform backdoor injection into the encoder component of the shadow object detector, using an additional shadow dataset distinct from the downstream dataset. The method for generating the backdoor data is consistent with the previous description in NA. Specifically, we add a trigger pattern to all target bboxes of the attacked class in the images of the benign dataset $D_{clean}$ and modify the corresponding annotations’ true labels to the target label $c_{t}$, resulting in the shadow dataset $D_{shadow}$:
$\mathcal{D}_{shadow}=\left\{\left(\boldsymbol{x}_{poisoned},\boldsymbol{y}_{target})\right)\mid(\boldsymbol{x},\boldsymbol{y})\in\mathcal{D}_{shadow}\right\}$ (4)
Once the shadow dataset $D_{shadow}$ is generated according to the above method, we adopt it to backdoor train with the standard training procedure, i.e.:
$\min\frac{1}{N}\sum_{(\boldsymbol{x},\boldsymbol{y})\in\mathcal{D}_{shadow}}\mathcal{L}\left(f(\boldsymbol{x}),\boldsymbol{y}\right)$ (5)
where $\mathcal{L}$ is the loss function, such as cross-entropy.
During training, the model parameters are adjusted by optimizing the loss function, typically achieved through stochastic gradient descent [20] and back-propagation. During this process, DNNs learn the mapping between trigger patterns and target labels. Then the attacker extracts the backdoor encoder from the shadow object detector obtained through backdoor training, and releases it.
Poisoning Fine-tuning. The victim obtains and uses the backdoored encoder from the previous step to construct their downstream detector, which inherits the encoder’s backdoor behavior. Additionally, the attacker embeds triggers into a portion of the training data and modifies the labels of these samples to the attacker’s predetermined target labels, creating and releasing a poisoned dataset. Notably, the poisoning proportion in the dataset is extremely small (e.g., 0.5%), making the attack more realistic. When the victim collects data from the network as a fine-tuning dataset, it may include the attacker’s poisoned data. The victim then uses this data to fine-tune their downstream tasks, following the process described in NA.
@@
TABLE I: Attack Performance (%) of Two Attacks on Object Detection. Note that $\uparrow$”/$\downarrow$”/“-” indicate the metric should be high/low/similar to the same metric of $M_{benign}$ to show the success of the attack.
Downstream Dataset VOC2007 MSCOCO
Model Fast R-CNN Fast R-CNN Mask R-CNN
Attack Method NA DSBA NA DSBA NA DSBA
$mAP_{benign}$ 68.69 70.75 51.76 53.03 53.45 53.97
$AP_{benign}$ 70.26 68.19 50.54 52.58 60.45 55.57
$mAP_{attack}$uparrow 62.77 66.25 48.64 50.27 50.26 51.04
$AP_{attack}$uparrow 80.01 89.48 45.54 55.98 49.52 56.44
$mAP_{attack+benign}$ 60.47 62.84 47.19 48.72 48.79 49.47
$AP_{attack+benign}$downarrow 1.46 1.54 0.87 0.0 0.0 0.0
$-mAP_{part\_attack+benign}\text{-}$ 58.98 60.54 46.12 47.54 44.77 48.42
ASR $\uparrow$ 72.56 86.55 69.95 70.33 70.51 71.63
IV Evaluation
IV-A Experiment settings
Datasets. In this work, we utilize the PASCAL VOC2007 [18], combined PASCAL VOC07+12 [21], and MSCOCO [22] datasets for comprehensive evaluations. Each image is annotated with bboxes coordinates and corresponding classes. The VOC2007 dataset consists of 20 categories, split into training/validation data (trainval) and testing data (test). The MSCOCO trainval dataset includes 80 categories, with 118k training images used for training purposes (58k for fine-tuning and 60k for shadow datasets), while the validation set serves as the testing data. In all experiments, we select the ”person” category as the attacked class and the ”dog” category as the target class.